QUESTION 1

  1. A 105 g spring-driven toy car is on circular track (radius 9.01 m) moving 30.7 m/s when the spring runs down. If, after 8.37 s, friction has slowed the car down to 4.03 m/s: findF, the magnitude of the net force acting on the car at that moment, in N.

    QUESTION 2

  2. A massless vertical rope is used to lower a block of mass 99.7 kg at constant acceleration5.09 m/s2 downward.  Find the work done by the tension in the rope, in J, after the block is lowered a distance of 44 m.
Answer #1

1.

Calculate net acceleration

a=(vf-vi)/t=(4.03-30.7)/8.37=3.2m/s^2

Calculate net force

Fnet= ma=0.105kg*3.2m/s^2=0.336N